The length of a room is 4 m more than 3 times its width. If the perimeter is 28m, how wide is the room?
Answer by Cromlix(4381)

Width = x
Length = 4 + 3x
Perimeter = 2*width + 2*length
28 = 2(x) + 2(4 + 3x)
28 = 2x + 8 + 6x
28 - 8 = 8x
20 = 8x
x = 20/8
x = 2.5m
Width = 2.5m
Hope this helps.
:-)